# Adaptive sampling accelerates the hybrid deviational particle simulations

### v1

@@ Abstract
To avoid ineffective collisions between the equilibrium states, the hybrid method with deviational particles (HDP) has been proposed to integrate the Fokker-Planck-Landau system, while leaving a new issue in sampling deviational particles from the high-dimensional source term. In this paper, we present an adaptive sampling (AS) strategy that first adaptively reconstructs a piecewise constant approximation of the source term based on sequential clustering via discrepancy estimation, and then samples deviational particles directly from the resulting adaptive piecewise constant function without rejection. The mixture discrepancy, which can be easily calculated thanks to its explicit analytical expression, is employed as a measure of uniformity instead of the star discrepancy the calculation of which is NP-hard. The resulting method, dubbed the HDP-AS method, runs approximately ten times faster than the HDP method while keeping the same accuracy in the Landau damping, two stream instability, bump on tail and Rosenbluth’s test problem.
@@
Fokker-Planck-Landau system; Deviational particles; Adaptive sampling; Piecewise constant reconstruction; Discrepancy
@@ 1 Introduction
As a fundamental mathematical model for the kinetic theory, the Fokker-Planck-Landau (FPL) system
$\partial_{t}f+\boldsymbol{v}\cdot\nabla_{\boldsymbol{x}}f+\boldsymbol{E}\cdot\nabla_{\boldsymbol{v}}f$ $=Q(f,f),$ (1)
@@
describes the evolution of the charged particles in phase space $(\boldsymbol{x},\boldsymbol{v})\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ [1, 2], where $f=f(\boldsymbol{x},\boldsymbol{v},t)$ gives the time-dependent distribution function, $\rho(\bm{x},t)=\int_{\mathbb{R}^{3}}f(\boldsymbol{x},\boldsymbol{v},t)\textup{d}\bm{v}$ the spatial density and $Q(f,f)$ the Landau collision operator. Numerical resolution of the FPL system encounters some fundamental obstacles due to the notorious curse of dimensionality and thus particle-based methods play a prevalent role benefiting from their (possible) dimension independence. The DSMC-PIC method, an often-used particle method, which exploits the operator splitting technique and combines the direct simulation Monte Carlo (DSMC) (such as the TA method by Takizuka and Abe [3] and the Nanbu method [4]) for the collision step
@@
and the particle in cell (PIC) [5, 6] for the transport step
$\partial_{t}f+\boldsymbol{v}\cdot\nabla_{\boldsymbol{x}}f+\boldsymbol{E}\cdot\nabla_{\boldsymbol{v}}f$ $=0,$ (3)
@@
e.g., the micro-macro approach for the BGK operator [7], the thermalization/dethermalization method for the spatial homogeneous Coulomb collisions [8] and the low-variance deviational simulation Monte Carlo methods for rarefied gas [9, 10]. For the FPL system (1), a hybrid method with deviational particles (HDP) was proposed along this line and allows the deviation $f_{d}$ to be negative [11, 12]. Namely, the HDP method decomposes the distribution as follows
@@
We propose an HDP method based on adaptive sampling (AS) for solving the FPL system (1). The resulting method, dubbed the HDP-AS method, reconstructs an adaptive piecewise constant approximation of $Q(f_{p}-f_{n},m)$ from the point distribution, which can be readily sampled from (see Figure 2 for a quick demonstration), and the wall time consumed by sampling new deviational particles is significantly reduced (see Figure 1). It partially borrows the idea from the discrepancy-based density estimation [15], which employs the star discrepancy to measure irregularity of points distribution and then constructs an adaptive approximation of high-dimensional unknown function. Considering the fact that the complexity of calculating the star discrepancy is NP-hard [16] and a high computational cost may be inevitable for a large sample sizes, HDP-AS replaces the star discrepancy with the mixture discrepancy which can be calculated explicitly [17]. Numerical results for the Landau damping, two stream instability, bump on tail and Rosenbluth’s test problem demonstrates that HDP-AS achieves about ten times acceleration compared with HDP while maintaining the same accuracy.
The rest of this paper is organized as follows. In Section 2, we first present some preliminaries of the FPL system, and then introduce related numerical attempts, including the TA method and the HDP method, which lay the groundwork for the proposed HDP-AS method. Section 3 details the HDP-AS method. Numerical experiments are conducted in Section 4 for comparing HDP-AS with HDP. The paper is concluded in Section 5.
@@ 2 Preliminaries and background
A hybrid representation of the distribution function, which involves both grids and particles, is widely used to integrate the FPL system (1). In the spatial direction, uniform grids $\mathcal{C}_{k},\,k=1,\dots,N_{x}$ are constructed with $\bm{x}_{k}$ being the center point of $\mathcal{C}_{k}$. The macroscopic quantities $\rho$, $\boldsymbol{u}$ and $T$ depend only on the position, and thus are defined on the spatial grids. In the velocity direction, the particle representation is employed. The positive part $f_{p}$ and negative part $f_{n}$ are represented by particles called P-particles and N-particles respectively. Denote the sets of these particles at $\bm{x}_{k}$ and time $t$ as $\mathbb{P}_{k}(t)$ and $\mathbb{N}_{k}(t)$, the size of which is $N_{p,k}$ and $N_{n,k}$, respectively. Below we illustrate how the TA [3] and HDP [12] methods simulate collisions in one time step $\Delta t$, which lay the groundwork for the proposed HDP-AS method.
@@ 2.1 The TA method
where $\Delta\boldsymbol{u}=(\bm{v}^{h}-\bm{v}^{g})-(\widetilde{\bm{v}}^{h}-\widetilde{\bm{v}}^{g})$ is the change of relative velocity. During the collision process, the magnitude of the relative velocity remains unchanged, but its direction is altered by the scattering angle $\Theta$ and azimuthal angle $\Phi$, which characterize the nature of the Coulomb collisions. In the TA method, $\Phi$ is chosen randomly with a uniform distribution in $\left[0,2\pi\right]$ and $\Theta=2\arctan(\delta)$, where $\delta$ is sampled from the normal distribution $\mathcal{N}(0,C\Delta t)$ and $C$ is determined by some physical constants as well as the magnitude of the relative velocity (see [3] for more details). Let $\bm{v}^{h}-\bm{v}^{g}:=(u_{x},u_{y},u_{z})$ and $u:=\sqrt{u_{x}^{2}+u_{y}^{2}+u_{z}^{2}}$. The change of relative velocity $\Delta\boldsymbol{u}:=(\Delta u_{x},\Delta u_{y},\Delta u_{z})$ can be computed by
@@
Algorithm 1 Simulate the G-H collisions via TA.
@@ 2.2 The HDP method
explicitly removing the collisions $Q(m,m)$ of the equilibrium parts, where $h_{\pm}=\frac{1}{2}\left(|h|\pm h\right)$ gives the positive/negative part. Here the notation $\tilde{f}$ is used to distinguish it from the desired solution $f=m+f_{p}-f_{n}$ in Eq. (4). It is introduced in HDP as a coarse approximation of $f$ and used in simulating collisions $Q(\tilde{f},f_{p})$ and $Q(\tilde{f},f_{n})$. The distribution $\tilde{f}$ is represented by F-particles and the set of these particles at $\bm{x}_{k}$ and time $t$ is denoted as $\mathbb{F}_{k}(t)$, the size of which is $N_{f,k}$. A key ingredient is that F-particles are not necessary to be a very accurate approximation of $f$, and one only needs $N_{f,k}\geq N_{p,k}+N_{n,k}$ to ensure there are enough F-particles to simulate collisions $Q(\tilde{f},f_{p})$ and $Q(\tilde{f},f_{n})$. An error analysis shows that F-particles introduces a negligible error (see Section 3.3 in [11]). Eq. (11) guides the evolution of the F-particles, P-particles and N-particles in a prediction–correction manner. The F-particles act as a prediction of $f$, and they are still simulated by the TA method. As corrections, P-particles and N-particles contain two parts in their evolution: One is collisions with F particles, simulated using the TA method; the other is sampling new particles from the source term
@@
Algorithm 2 Simulate the Coulomb collisions in the HDP method.
@@ 3 Adaptive sampling
Different from the acceptance-rejection sampling in the HDP method [11, 12], HDP-AS proposes an adaptive high-dimensional sampling approach to sample from $\Delta m_{k}(\bm{v})_{\pm}$ required by Algorithm 2. There exist two steps in the HDP-AS method: (a) Provide an adaptive piecewise constant approximation of the function $\Delta m_{k}(\bm{v})$, denoted by $\overline{\Delta m_{k}}(\bm{v})$, and (b) sample new deviational particles directly from $\overline{\Delta m_{k}}(\bm{v})$ instead of $\Delta m_{k}(\bm{v})$. In contrast to the acceptance-rejection sampling, the implementation of (b) is much more straightforward and without rejection. In the following, we only detail how to achieve (a) in an efficient manner.
It involves two processes: (a1) Simulate the Coulomb collisions $Q(f_{p},m)$ and $Q(f_{n},m)$ in the function $\Delta m_{k}(\bm{v})$ via the TA method in Algorithm 1 and obtain the point distribution $W_{k}(\bm{v})$ (see Eq. (13)), approximating $Q\left(f_{p},m\right)-Q\left(f_{n},m\right)$ in the weak sense, and (a2) adaptively reconstruct $\overline{\Delta m_{k}}(\bm{v})$ from the point distribution $W_{k}(\bm{v})$. Denote the particles with the equilibrium distribution $m$ as M-particles. In $\Delta m_{k}(\bm{v})$, $Q(f_{p},m)$ and $Q(f_{n},m)$ represent that P-M collisions and N-M collisions respectively, both of which can be simulated by the TA method. Let $\mathbb{M}^{+}_{k}$ and $\mathbb{M}^{-}_{k}$ collect all the M-particles after P-M collisions and N-M collisions at grid point $\bm{x}_{k}$ and time $t$, respectively. They can form a point distribution
$W_{k}(\bm{v})=\frac{\rho_{p}(\bm{x}_{k},t)}{N_{p,k}}\sum_{\boldsymbol{a}\in\mathbb{M}^{+}_{k}}\delta(\bm{v}-\boldsymbol{a})-\frac{\rho_{n}(\bm{x}_{k},t)}{N_{n,k}}\sum_{\boldsymbol{a}\in\mathbb{M}^{-}_{k}}\delta(\bm{v}-\boldsymbol{a}),$ (13)
where
$\rho_{p}(\bm{x},t)=\int_{\mathbb{R}^{3}}f_{p}(\boldsymbol{x},\boldsymbol{v},t)\textup{d}\bm{v},\quad\rho_{n}(\bm{x},t)=\int_{\mathbb{R}^{3}}f_{n}(\boldsymbol{x},\boldsymbol{v},t)\textup{d}\bm{v},$ (14)
and $\delta(\bm{v})$ is the Dirac distribution in $\mathbb{R}^{3}$. Actually, the TA method ensures that $W_{k}$ in Eq. (13) may approximate $Q\left(f_{p},m\right)-Q\left(f_{n},m\right)$ in the weak sense, i.e., $\forall\,\varphi(\boldsymbol{v})\in\mathbb{H}_{0}^{2}\left(\mathbb{R}^{3}\right)$, $\forall\,k\in\{1,\dots,N_{x}\}$,
$\left\langle\varphi(\bm{v}),Q\left(f_{p}(\bm{x}_{k},\bm{v},t),m(\bm{x}_{k},\bm{v},t)\right)-Q\left(f_{n}(\bm{x}_{k},\bm{v},t),m(\bm{x}_{k},\bm{v},t)\right)\right\rangle\approx\left\langle\varphi(\bm{v}),W_{k}(\bm{v})\right\rangle$ (15)
@@
$\left\langle\varphi(\bm{v}),W_{k}(\bm{v})\right\rangle=\frac{\rho_{p}(\bm{x}_{k},t)}{N_{p,k}}\sum_{\boldsymbol{a}\in\mathbb{M}^{+}_{k}}\varphi\left(\boldsymbol{a}\right)-\frac{\rho_{n}(\bm{x}_{k},t)}{N_{n,k}}\sum_{\boldsymbol{a}\in\mathbb{M}^{-}_{k}}\varphi\left(\boldsymbol{a}\right),$ (16)
where the standard inner product $\langle g,h\rangle=\int_{\mathbb{R}^{3}}g(\bm{v})h(\bm{v})\textup{d}\bm{v}$ is adopted. Algorithm 3 details the way of obtaining $\mathbb{M}^{\pm}_{k}$ in Eq. (13).
Algorithm 3 Simulate the Coulomb collisions in $\Delta m_{k}(\bm{v})$ for $k=1,\dots,N_{x}$.
Input: Grid number $N_{x}$, the set of P-particles $\mathbb{P}_{k}(t)$, the set of N-particles $\mathbb{N}_{k}(t)$, the macroscopic velocity $\boldsymbol{u}(\bm{x}_{k},t)$ and temperature $T(\bm{x}_{k},t)$, $k=1,\dots,N_{x}$.
Output: Particle sets $\mathbb{M}^{\pm}_{k}$ in Eq. (13), $k=1,\dots,N_{x}$.
@@
$\approx\frac{\rho_{p}(\bm{x}_{k},t)}{N_{p,k}}\sum_{\boldsymbol{a}\in\mathbb{M}^{+}_{k}}\frac{\mathbbm{1}_{\Omega_{l}}\left(\boldsymbol{a}\right)}{\mu\left(\Omega_{l}\right)}-\frac{\rho_{n}(\bm{x}_{k},t)}{N_{n,k}}\sum_{\boldsymbol{a}\in\mathbb{M}^{-}_{k}}\frac{\mathbbm{1}_{\Omega_{l}}\left(\boldsymbol{a}\right)}{\mu\left(\Omega_{l}\right)},$
@@
Algorithm 4 Reconstruct the adaptive piecewise constant approximation $\overline{\Delta m_{k}}(\bm{v})$.
@@
$L=1,\Omega_{1}=\Omega,\mathcal{P}_{1}=\left\{\Omega_{1}\right\}$;
2:$N_{p,k}\leftarrow$ the size of $\mathbb{M}^{+}_{k}$;
3:$N_{n,k}\leftarrow$ the size of $\mathbb{M}^{-}_{k}$;
4:$N=N_{p,k}+N_{n,k}$;
@@
12:         Calculate the discrepancies $D(\mathbb{M}^{+}_{\text{cur}})$ and $D(\mathbb{M}^{-}_{\text{cur}})$;
@@
16:              $\Omega_{l}\leftarrow\Omega_{l}^{(1)},\Omega_{L+1}\leftarrow\Omega_{l}^{(2)}$;
@@
28:Calculate piecewise constants $c_{l},\,l=1,\dots,L$ according to Eq. (19);
29:$\overline{\Delta m_{k}}(\bm{v})\leftarrow\sum\limits_{l=1}^{L}c_{l}\Delta t\mathbbm{1}_{\Omega_{l}}(\bm{v})$;
@@
    Where to split: Suppose $\Omega_{l}=\left[a_{1}^{(l)},b_{1}^{(l)}\right]\times\left[a_{2}^{(l)},b_{2}^{(l)}\right]\times\left[a_{3}^{(l)},b_{3}^{(l)}\right]$ and $s_{i_{0},j_{0}}^{(l)}=a_{j_{0}}^{(l)}+\frac{i_{0}}{q}\left(b_{j_{0}}^{(l)}-a_{j_{0}}^{(l)}\right)$ are some partition points to be selected where $i_{0}=1,2,\ldots,q-1$, $j_{0}=1,2,3$. To further divide the non-uniform sub-domains, the (sub)-optimal split points, denoted by $\widetilde{s}_{i_{0},j_{0}}^{(l)}$, are suggested to maximize the difference gap
@@
    Whether to split: We use the discrepancies $D(\mathbb{M}^{+}_{\text{cur}})$ and $D(\mathbb{M}^{-}_{\text{cur}})$ as indicators to measure the uniformity of the points distribution. Small discrepancies indicate a relatively uniform particle distribution and no further partitioning is necessary. Discrepancy is a broad category and contains various specific definitions [18, 19]. The star discrepancy is used in [15], the definition of which for point set $\left(\bm{v}_{1},\dots,\bm{v}_{n}\right)\subset\left[0,1\right]^{3n}$ is stated as follows
@@
    $D^{\text{mix}}(\mathbb{M}^{+}_{\text{cur}})$ $=D^{\text{mix}}(\iota_{l}(\bm{v}_{1}^{+}),\dots,\iota_{l}(\bm{v}_{N_{+}}^{+})),$ (24)
@@
    where the parameter $\vartheta$ governs the depth of partition, and $N,N_{\pm}$ are given in Lines 4, 10 and 11 in Algorithm 4. A smaller $\vartheta$ leads to a finer partitioning.
In contrast to the low acceptance rate of the acceptance-rejection method employed by HDP (see Figure 4), sampling from the piecewise constant functions $\overline{\Delta m_{k}}(\bm{v})$ in HDP-AS is straightforward and without rejection, thereby significantly enhancing the sampling efficiency. Using this improvement, the main time-consuming steps of HDP-AS become simulating collisions and transport, while sampling new deviational particles is no longer time-consuming now (see Figure 1).

### v2

@@ Abstract
To avoid ineffective collisions between the equilibrium states, the hybrid method with deviational particles (HDP) has been proposed to integrate the Vlasov-Poisson-Landau system, while leaving a new issue in sampling deviational particles from the high-dimensional source term. In this paper, we present an adaptive sampling (AS) strategy that first adaptively reconstructs a piecewise constant approximation of the source term based on sequential clustering via discrepancy estimation, and then samples deviational particles directly from the resulting adaptive piecewise constant function without rejection. The mixture discrepancy, which can be easily calculated thanks to its explicit analytical expression, is employed as a measure of uniformity instead of the star discrepancy the calculation of which is NP-hard. The resulting method, dubbed the HDP-AS method, samples deviational particles through adaptive sampling instead of the acceptance-rejection method in the original HDP method. In the Landau damping, two stream instability, bump on tail and Rosenbluth’s test problems, the HDP-AS method runs approximately ten times faster than the HDP method while keeping the same accuracy.
@@
Vlasov–Poisson–Landau system; Deviational particles; Adaptive sampling; Piecewise constant reconstruction; Discrepancy
@@ 1 Introduction
As a fundamental mathematical model for the kinetic theory, the Vlasov–Poisson–Landau (VPL) system
$\partial_{t}f+\boldsymbol{v}\cdot\nabla_{\boldsymbol{x}}f-\boldsymbol{E}\cdot\nabla_{\boldsymbol{v}}f$ $=Q(f,f),$ (1)
@@
describes the evolution of the charged particles in phase space $(\boldsymbol{x},\boldsymbol{v})\in\mathbb{R}^{3}\times\mathbb{R}^{3}$ [1, 2], where $f=f(\boldsymbol{x},\boldsymbol{v},t)$ gives the time-dependent distribution function, $\rho(\bm{x},t)=\int_{\mathbb{R}^{3}}f(\boldsymbol{x},\boldsymbol{v},t)\textup{d}\bm{v}$ the spatial density and $Q(f,f)$ the Landau collision operator. The form of the VPL system (1) adopted in this paper is consistent with that presented in [3]. Numerical resolution of the VPL system encounters some fundamental obstacles due to the notorious curse of dimensionality and thus particle-based methods play a prevalent role benefiting from their (possible) dimension independence. The DSMC-PIC method, an often-used particle method, which exploits the operator splitting technique and combines the direct simulation Monte Carlo (DSMC) (such as the TA method by Takizuka and Abe [4] and the Nanbu method [5]) for the collision step
@@
and the particle in cell (PIC) [6, 7] for the advection step
$\partial_{t}f+\boldsymbol{v}\cdot\nabla_{\boldsymbol{x}}f-\boldsymbol{E}\cdot\nabla_{\boldsymbol{v}}f$ $=0,$ (3)
@@
e.g., the micro-macro approach for the BGK operator [8], the thermalization/dethermalization method for the spatial homogeneous Coulomb collisions [9] and the low-variance deviational simulation Monte Carlo methods for rarefied gas [10, 11]. For the VPL system (1), a hybrid method with deviational particles (HDP) was proposed along this line and allows the deviation $f_{d}$ to be negative [12, 3]. Namely, the HDP method decomposes the distribution as follows
@@
We propose an HDP method based on adaptive sampling (AS) for solving the VPL system (1). The resulting method, dubbed the HDP-AS method, reconstructs an adaptive piecewise constant approximation of $Q(f_{p}-f_{n},m)$ from the point distribution, which can be readily sampled from (see Figure 2 for a quick demonstration), and the wall time consumed by sampling new deviational particles is significantly reduced (see Figure 1). This two-phase architecture — first constructing a piecewise constant approximation of the target distribution and then sampling from the resulting piecewise constant approximation — can trace its lineage to the 1960s ziggurat algorithm [15, 16] and dramatically enhance the sampling efficiency. Moreover, HDP-AS partially borrows the idea from the discrepancy-based density estimation [17], which employs the star discrepancy to measure irregularity of points distribution and then constructs an adaptive approximation of high-dimensional unknown function. Considering the fact that the complexity of calculating the star discrepancy is NP-hard [18] and a high computational cost may be inevitable for a large sample sizes, HDP-AS replaces the star discrepancy with the mixture discrepancy which can be calculated explicitly [19]. The only difference between HDP and HDP-AS lies in that HDP samples new deviational particles by the acceptance-rejection method, whereas HDP-AS employs adaptive sampling. The other steps remain the same for both methods. Numerical results for the Landau damping, two stream instability, bump on tail and Rosenbluth’s test problem demonstrates that HDP-AS achieves about ten times acceleration compared with HDP while maintaining the same accuracy.
The rest of this paper is organized as follows. In Section 2, we first present some preliminaries of the VPL system, and then introduce related numerical attempts, including the TA method and the HDP method, which lay the groundwork for the proposed HDP-AS method. Section 3 details the HDP-AS method. Numerical experiments are conducted in Section 4 for comparing HDP-AS with HDP. The paper is concluded in Section 5. This paper addresses sampling new deviational particles from the source term in the collision step of the HDP method. Other components like the advection step’s macro-micro decomposition and collision step’s particle annihilation remain identical between HDP-AS and HDP and are thus omitted here for brevity. The interested readers can refer to [12, 3] for more details.
@@ 2 Preliminaries and background
A hybrid representation of the distribution function, which involves both grids and particles, is widely used to integrate the VPL system (1). In the spatial direction, uniform grids $\mathcal{C}_{k},\,k=1,\dots,N_{x}$ are constructed with $\bm{x}_{k}$ being the center point of $\mathcal{C}_{k}$. The macroscopic quantities $\rho$, $\boldsymbol{u}$ and $T$ depend only on the position, and thus are defined on the spatial grids. In the velocity direction, the particle representation is employed. The positive part $f_{p}$ and negative part $f_{n}$ are represented by particles called P-particles and N-particles respectively. Denote the sets of these particles at $\bm{x}_{k}$ and time $t$ as $\mathbb{P}_{k}(t)$ and $\mathbb{N}_{k}(t)$, the size of which is $N_{p,k}$ and $N_{n,k}$, respectively.
To solve the VPL system (1), we also start with an operator splitting between the collision part (2) and the advection part (3). In the HDP-AS method, the simulation of advection part employs the identical micro-macro decomposition framework [20] as that utilized in HDP, so it will not be repeated here. Below we illustrate how the TA [4] and HDP [3] methods simulate collisions in one time step $\Delta t$, which lay the groundwork for the proposed HDP-AS method.
@@ 2.1 The TA method
where $\Delta\boldsymbol{u}=(\widetilde{\bm{v}}^{h}-\widetilde{\bm{v}}^{g})-(\bm{v}^{h}-\bm{v}^{g})$ is the change of relative velocity. During the collision process, the magnitude of the relative velocity remains unchanged, but its direction is altered by the scattering angle $\Theta$ and azimuthal angle $\Phi$, which characterize the nature of the Coulomb collisions. In the TA method, $\Phi$ is chosen randomly with a uniform distribution in $\left[0,2\pi\right]$ and $\Theta=2\arctan(\delta)$, where $\delta$ is sampled from the normal distribution $\mathcal{N}(0,C\Delta t)$ and $C$ is determined by some physical constants as well as the magnitude of the relative velocity (see [4] for more details). Let $\bm{v}^{h}-\bm{v}^{g}:=(u_{x},u_{y},u_{z})$ and $u:=\sqrt{u_{x}^{2}+u_{y}^{2}+u_{z}^{2}}$. The change of relative velocity $\Delta\boldsymbol{u}:=(\Delta u_{x},\Delta u_{y},\Delta u_{z})$ can be computed by
@@
Algorithm 1 Simulate the G-H collisions with the TA method
@@ 2.2 The HDP method
explicitly removing the collisions $Q(m,m)$ of the equilibrium parts, where $h_{\pm}=\frac{1}{2}\left(|h|\pm h\right)$ gives the positive/negative part. Here the notation $\tilde{f}$ is used to distinguish it from the desired solution $f=m+f_{p}-f_{n}$ in Eq. (4). It is introduced in HDP as a coarse approximation of $f$ and used in simulating collisions $Q(\tilde{f},f_{p})$ and $Q(\tilde{f},f_{n})$. The distribution $\tilde{f}$ is represented by F-particles and the set of these particles at $\bm{x}_{k}$ and time $t$ is denoted as $\mathbb{F}_{k}(t)$, the size of which is $N_{f,k}$. The three parts $\tilde{f},f_{p},f_{n}$ are presented respectively by the F-particles, P-particles and N-particles, i.e., $\forall\,\varphi(\boldsymbol{v})\in\mathbb{H}_{0}^{2}\left(\mathbb{R}^{3}\right)$, $\forall\,k\in\{1,\dots,N_{x}\}$,
[…]
$\left\langle\varphi,f_{p}(\bm{x}_{k},\bm{v},t)\right\rangle$ $\approx N_{\text{eff }}\sum_{\bm{v}_{f_{p}}\in\mathbb{P}_{k}(t)}\varphi(\bm{v}_{f_{p}}),$
$\left\langle\varphi,f_{n}(\bm{x}_{k},\bm{v},t)\right\rangle$ $\approx N_{\text{eff }}\sum_{\bm{v}_{f_{n}}\in\mathbb{N}_{k}(t)}\varphi(\bm{v}_{f_{n}}),$
where the standard inner product $\langle g,h\rangle=\int_{\mathbb{R}^{3}}gh\textup{d}\bm{v}$ is adopted. The effective number $N_{\text{eff }}^{F}$ and $N_{\text{eff }}$ represent the physical densities per numerical particle and satisfy
$N_{f,k}N_{\text{eff }}^{F}\approx\rho(\bm{x}_{k},t)=\int_{\mathbb{R}^{3}}\tilde{f}(\bm{x}_{k},\bm{v},t)\textup{d}\bm{v},$ (13)
$N_{p,k}N_{\text{eff }}\approx\rho_{p}(\bm{x}_{k},t)=\int_{\mathbb{R}^{3}}f_{p}(\bm{x}_{k},\bm{v},t)\textup{d}\bm{v},$
$N_{n,k}N_{\text{eff }}\approx\rho_{n}(\bm{x}_{k},t)=\int_{\mathbb{R}^{3}}f_{n}(\bm{x}_{k},\bm{v},t)\textup{d}\bm{v}.$
Eq. (13) can be obtained by taking $\varphi\equiv 1$ in Eq. (12). A key ingredient is that F-particles are not necessary to be a very accurate approximation of $f$, and one only needs $N_{f,k}\geq N_{p,k}+N_{n,k}$ to ensure there are enough F-particles to simulate collisions $Q(\tilde{f},f_{p})$ and $Q(\tilde{f},f_{n})$. An error analysis shows that F-particles introduces a negligible error (see Section 3.3 in [12]). Eq. (11) guides the evolution of the F-particles, P-particles and N-particles in a prediction–correction manner. The F-particles act as a prediction of $f$, and they are still simulated by the TA method. As corrections, P-particles and N-particles contain two parts in their evolution: One is collisions with F particles, simulated using the TA method; the other is sampling new particles from the source term
@@
Algorithm 2 Simulate the Coulomb collisions with the HDP method
@@ 3 Adaptive sampling
Different from the acceptance-rejection sampling in the HDP method [12, 3], HDP-AS proposes an adaptive high-dimensional sampling approach to sample from $\Delta m_{k}(\bm{v})_{\pm}$ required by Algorithm 2. There exist two steps in the HDP-AS method: (a) Provide an adaptive piecewise constant approximation of the function $\Delta m_{k}(\bm{v})$, denoted by $\overline{\Delta m_{k}}(\bm{v})$, and (b) sample new deviational particles directly from $\overline{\Delta m_{k}}(\bm{v})$ instead of $\Delta m_{k}(\bm{v})$. In contrast to the acceptance-rejection sampling, the implementation of sampling from a piecewise constant function in (b) is much more straightforward and without rejection. Moreover, we employ adaptive piecewise constants instead of uniform ones to alleviate the curse of dimensionality. As the implementation of (b) is straightforward, we focus exclusively on how to achieve (a) in an efficient manner in the following.
It involves two processes: (a1) Simulate the Coulomb collisions $Q(f_{p},m)$ and $Q(f_{n},m)$ in the function $\Delta m_{k}(\bm{v})$ via the TA method in Algorithm 1 and obtain the point distribution $W_{k}(\bm{v})$ (see Eq. (15)), approximating $Q\left(f_{p},m\right)-Q\left(f_{n},m\right)$ in the weak sense, and (a2) adaptively reconstruct $\overline{\Delta m_{k}}(\bm{v})$ from the point distribution $W_{k}(\bm{v})$. Denote the particles with the equilibrium distribution $m$ as M-particles, which are not tracked and only sampled from (5) whenever needed. In $\Delta m_{k}(\bm{v})$, $Q(f_{p},m)$ and $Q(f_{n},m)$ represent that P-M collisions and N-M collisions respectively, both of which can be simulated by the TA method. Let $\mathbb{M}^{+}_{k}$ and $\mathbb{M}^{-}_{k}$ collect all the M-particles after P-M collisions and N-M collisions at grid point $\bm{x}_{k}$ and time $t$, respectively. They can form a point distribution
$W_{k}(\bm{v})=N_{\text{eff }}\sum_{\boldsymbol{a}\in\mathbb{M}^{+}_{k}}\delta(\bm{v}-\boldsymbol{a})-N_{\text{eff }}\sum_{\boldsymbol{a}\in\mathbb{M}^{-}_{k}}\delta(\bm{v}-\boldsymbol{a}),$ (15)
where $\delta(\bm{v})$ is the Dirac distribution in $\mathbb{R}^{3}$. Actually, the TA method ensures that $W_{k}$ in Eq. (15) may approximate $Q\left(f_{p},m\right)-Q\left(f_{n},m\right)$ in the weak sense, i.e., $\forall\,\varphi(\boldsymbol{v})\in\mathbb{H}_{0}^{2}\left(\mathbb{R}^{3}\right)$, $\forall\,k\in\{1,\dots,N_{x}\}$,
$\left\langle\varphi,Q\left(f_{p}(\bm{x}_{k},\bm{v},t),m(\bm{x}_{k},\bm{v},t)\right)-Q\left(f_{n}(\bm{x}_{k},\bm{v},t),m(\bm{x}_{k},\bm{v},t)\right)\right\rangle\approx\left\langle\varphi,W_{k}\right\rangle$ (16)
@@
$\left\langle\varphi,W_{k}\right\rangle=N_{\text{eff }}\sum_{\boldsymbol{a}\in\mathbb{M}^{+}_{k}}\varphi\left(\boldsymbol{a}\right)-N_{\text{eff }}\sum_{\boldsymbol{a}\in\mathbb{M}^{-}_{k}}\varphi\left(\boldsymbol{a}\right).$ (17)
Algorithm 3 details the way of obtaining $\mathbb{M}^{\pm}_{k}$ in Eq. (15).
Algorithm 3 Simulate the Coulomb collisions in $\Delta m_{k}(\bm{v})$ for $k=1,\dots,N_{x}$
Input: Grid number $N_{x}$, the set of P-particles $\mathbb{P}_{k}(t)$, the set of N-particles $\mathbb{N}_{k}(t)$, the macroscopic velocity $\boldsymbol{u}(\bm{x}_{k},t)$ and temperature $T(\bm{x}_{k},t)$.
Output: Particle sets $\mathbb{M}^{\pm}_{k}$ in Eq. (15).
@@
[…]
@@
Algorithm 4 Reconstruct the adaptive piecewise constant approximation $\overline{\Delta m_{k}}(\bm{v})$
@@
[…]
@@
[…]
@@
16:     $\Omega_{l}\leftarrow\Omega_{l}^{(1)},Q_{L+1}\leftarrow\Omega_{l}^{(2)}$;
@@
28:Calculate piecewise constants $c_{l},\,l=1\dots,L$ according to Eq. (20);
[…]
@@
    Where to split: Suppose $\Omega_{l}=\left[a_{1}^{(l)},b_{1}^{(l)}\right]\times\left[a_{2}^{(l)},b_{2}^{(l)}\right]\times\left[a_{3}^{(l)},b_{3}^{(l)}\right]$ and $s_{i_{0},j_{0}}^{(l)}=a_{j_{0}}^{(l)}+\frac{i_{0}}{q}\left(b_{i_{0}}^{(l)}-a_{i_{0}}^{(l)}\right)$ are some partition points to be selected where $i_{0}=1,2,\ldots,q-1$, $j_{0}=1,2,3$. To further divide the non-uniform sub-domains, the (sub)-optimal split points, denoted by $\widetilde{s}_{i_{0},j_{0}}^{(l)}$, are suggested to maximize the difference gap
@@
[…]
@@
[…]
@@
    where the parameter $\vartheta$ governs the depth of partition, and $N,N_{\pm}$ are given in Lines 4, 10 and 11 in Algorithm 4. A smaller $\vartheta$ leads to a finer partitioning. The above criterion (27) treats the P-particles and N-particles on the same footing and has been successfully applied into an efficient stochastic particle simulation of the Wigner-Coulomb quantum dynamics [24].
In contrast to the low acceptance rate of the acceptance-rejection method employed by HDP (see Figure 4), sampling from the piecewise constant functions $\overline{\Delta m_{k}}(\bm{v})$ in HDP-AS is straightforward and without rejection, thereby significantly enhancing the sampling efficiency. Using this improvement, the main time-consuming steps of HDP-AS become simulating collisions and advection, while sampling new deviational particles is no longer time-consuming now (see Figure 1).
Remark 1.
The iconic ziggurat algorithm [15, 16] efficiently generates samples by partitioning the desired distribution functions into equal-area segments. Knuth praises it as “a very pretty example of mathematical theory intimately interwoven with programming ingenuity—a fine illustration of the art of computer programming!” in his famous book [25]. This work inherits the idea of the ziggurat algorithm to solve the sampling problem in the HDP method. The main challenges in current applications lie in the desired distribution $\Delta m_{k}(\bm{v})$ is a three-dimensional function and we need efficiently construct a piecewise constant approximation from its point cloud form (15). Our HDP-AS method addresses these specific challenges through an adaptive domain partitioning based on the mixture discrepancy.